OpenStudy (rational):

Fun fibonacci sequence problem: show that \[\dfrac{F_{n+1}}{F_n} =1+ \dfrac{1}{1+\dfrac{1}{1+\dfrac{1}{1+\dfrac{1}{\ddots \text{n times}{}}}}}\]

OpenStudy (rational):

\[1, 1, 2, 3, 5, 8, ... \] examples : \(\dfrac{F_{1+1}}{F_1} = 1\) \(\dfrac{F_{2+1}}{F_2} = 1+\dfrac{1}{1}\) \(\dfrac{F_{3+1}}{F_3} = 1+\dfrac{1}{1+\dfrac{1}{1}}\)

OpenStudy (dan815):